The csthesis Class for L^{A}TEX 2ε

Martin Reddy

September 20, 1996

1 Introduction

This document describes the csthesis class (v2.0) which has been specifically written for L^{A}TEX 2ε. This class provides a general style for presenting undergraduate and postgraduate theses in the department of Computer Science, which are in line with the official University regulations. To use this class, all you need do is begin your L^{A}TEX document with the declaration:

     \documentclass[<options>]{csthesis}

The default behaviour of this formatting class is to produce documents which are typeset as 12 point text on single-sided A4 pages with 1\frac{1}{2} line spacing. However, a number of options are provided by the csthesis class to alter these default formatting setting. Also, a collection of useful environments and commands are defined to simplify the production of your thesis.

This document will begin by introducing the various options which the csthesis class currently supports. It will also describe the various environments and commands which it offers the user. This is followed by an important note on printing theses within the department. Finally, an appendix is also provided which details the current University regulations for formatting theses; as well as a simple example of using the csthesis class.

2 Class Options

The csthesis class is built on top of the standard L^{A}TEX 2ε report class. As such, it will accept and support all of the options which can be passed to report, e.g. twoside, leqno, fleqn etc. In addition to these inherited features, csthesis provides the following options (which can be specified in the square-bracketed section of the \documentclass command.

phd : By supplying this option, the \maketitle command will produce a title page for a Ph.D. thesis (please refer to Section 4 for a description of the \maketitle command under the csthesis class). This is the default operation. mphil : Produces a title page for M.Phil. theses when the \maketitle command is used.

msc : Produces a title page for M.Sc. theses when the \maketitle command is used.

bsc : Produces a title page for B.Sc. theses when the \maketitle command is used.

deptreport : This option changes the layout of the title page which is produced by the \maketitle command so that the title and author are visible when the document is bound with one of the departmental (orange) card covers.

centerchapter : All chapter headings are centred on the page (this includes Appendicies, Table of Contents, Bibliography etc.). There is a synonym for this option called centrechapter. This is the default operation.

leftchapter : All chapter headings are positioned flush with the left hand margin (this includes Appendicies, Table of Contents, Bibliography etc.).

rightchapter : All chapter headings are positioned flush with the right hand margin (this includes Appendicies, Table of Contents, Bibliography etc.).

abbrevs : Exports a number of commands for common abbreviations (such as N.B. i.e. e.g. etc.). These are detailed in Section 4.

draft : The draft option will produce a single spaced thesis with smaller margins to reduce paper wastage. In addition to this, the standard draft behaviour of the report class will be implemented, e.g. all Encapsulated PostScript (EPS) figures will be replaced by a bounding box and a filename. The user may also wish to utilise the twoside option in conjuction with the -Zbook,duplex lpr options when printing a draft thesis in order to reduce paper wastage.

singlespacing : This option can be used to override the default 1\frac{1}{2} spacing behaviour and force a single spaced thesis.

doublespacing : This option can be used to override the default 1\frac{1}{2} spacing behaviour and force a double spaced thesis.

fullspacing : By default, L^{A}TEX uses a leading—the distance between the baseline of consecutive lines—which is 1.2 times the font size. I.e. for 12 point text, a single spaced document will have a leading of 14.4 points. The default definitions of oneand-a-half spacing and double spacing within the csthesis class are exactly 1.5 and 2.0 times the font size (i.e. 18 and 24 points respectively). If you prefer to include the additional 20% scaling, then the fullspacing option will incorporate this into the definition for one-and-a-half and double spacing (i.e. produce 28.8 and 21.6 point leading, respectively).

parskip : Alters the paragraph formatting so that each paragraph is separated by a vertical space, and there is no indentation at the start of each paragraph. (The list environments are also updated to reflect this change.)

usequotes : When using the quotation or citetext environments (see next section), the text is surrounded by quotation marks, e.g. “text”.

3 Environments

This section describes the various environments which are defined within the csthesis class. For example, if you wish to use an environment called foo, then you would enter:

\begin{foo}
  ... your text here ...
\end{foo}

abstract : As with the the report style, the csthesis class has an environment for you to enter your abstract into. This is similar to the standard report one, but with a couple of differences (e.g. the title is in a larger font and the abstract is not centred between the top and bottom margins).

acknowledgements : Normally, an Acknowledgements section follows the abstract. This can be created by simply using the acknowledgements environment.

declaration : You will be required to include a Declaration section to state that the thesis is your own work. This will normally follow the Acknowledgements section. Please also note the existence of the command, \standarddeclaration, described below in Section 4, which automatically produces a default Declaration page.

quotation : The University regulations (see Appendix A) state that quotations should be single spaced. The quotations environment has therefore been redefined to act in this fashion. It will also display the quotation in a slanted font. Using the usequotes class option, you can have quotation marks automatically included in the quotation environment.

citetext : This environment is very similar to the quotation environment, with the exception that it can accept a single parameter to define the name and details of the originator. This will then be right justified directly after the quotation. E.g.

    \begin{citetext}{Monty Python}
       And now for something completely different\ldots
    \end{citetext}

4 Commands

The csthesis class also exports a number of commands which can be used within your thesis. These are common or useful operations which are provided to ease the development of your L^{A}TEX 2ε document.

\maketitle : This command has been redefined for the csthesis class. Instead of the default report operation, it will produce a suitably formatted title page with the title of the thesis, the author, the degree, the university name and the year. For this, the user can define the following strings before calling the \maketitle command:

• \title{...} : the title of the thesis.

• \author{...} : the author of the thesis.

• \submityear{...} : the year of presentation.

• \graduationdate{...} : the date (Month Year) of the graduation ceremony (optional).

For example, the following extract illustrates how to create a thesis title page. (Recall that the phd (default), mphil, msc and bsc class options can be used to change the title of the degree which is listed on the title page.)

   \title{How I Did It}
   \author{Victor Frankenstein}
   \submityear{1785}
   \maketitle

\dedication : You can use this command to insert a short dedication page at the front of your thesis. This will usually follow immediately after the title page (i.e. before the abstract, acknowledgements and declaration). For example:

     \dedication{To my Mum.}

\standarddeclaration : As noted above, you will be expected to provide a Declaration page at the front of your thesis. You can use the declaration environment to construct your own such page if you wish. However, the \standarddeclaration command can be used to automatically create a Declaration page with a default declaration inserted.

\singlespace : This command sets the subsequent line spacing to be single spaced. For example, if you wanted to present your Bibliography in single spaced text, then you could enter the following:

    \singlespace
    \bibliography{<bibfilename>}

\oneandahalfspace : This command sets the subsequent line spacing to be 1.5 spaced. This is the default format used by the csthesis class.

\doublespace : This command sets the subsequent line spacing to be double spaced.

\thesiscaption[{<lof-caption>}]{<caption>}{<label>} : This command can be used in place of the L^{A}TEX \caption and \label commands in order to provide a consitent captioning for figures and tables within your thesis. It requires two parameters: a piece of text to be used as a caption, and a label to hold the figure number. The caption will be formatted in italics with a left and right margin slightly less than the main text margins. You may also specify an optional parameter in square brackets which will be used for the caption in any List of Figures/Tables section. This should only be used within figure, figure*, table or table* environments.

Abbreviations : The csthesis class also provides a number of simple commands for common abbreviations and symbols (N.B. you must specify the abbrevs class option in order to define these). These are described in the table below.^{1}

Command  Description  Example
\NB  nota bene (note well)  N.B.
\eg  exempli gratia (for example)  e.g.
\Eg  exempli gratia (for example)  E.g.
\ie  id est (in other words)  i.e.
\Ie  id est (in other words)  I.e.
\etc  et cetera (and other things)  etc.
\etal  et alii (and others)  et al.
\etseq  et sequens (and the following)  et al.
\role, \Role  accented version of role  rôle, Rôle
\naive, \Naive  accented version of naive  naïve, Naïve
\precis, \Precis accented version of precis  précis, Précis
\tm superscripted trademark symbol TM
\cpright  superscripted copyright symbol \begin{matrix}c \\ ◯ \end{matrix}
\degrees  degrees symbol  ∘

5 Important Notes on Printing

The csthesis class has been designed to produce theses with margins that correspond precisely with University regulations. However, this is confounded by the fact that different printers within the department print pages at a slightly different offset and scale. This discrepancy can be in the order of several millimetres between any two printers and so can be significant for a thesis. The default operation of csthesis is therefore to produce pages which conform to University regulations if they are printer on the hp24 laser printer.

Should you wish to print your thesis on another printer, then you can supply the command

\SetPrinter{<printername>} in your preamble (i.e. before the \begin{document}) to set

the correct margins for that printer. Currently supported values for <printername> are:

lw14, hp24, hp16 and mp16.

It should be noted that these customised margin settings are valid for the current configurations of the appropriate printers. However, these settings may change over time and new printers may be installed or old ones removed.

1Note that all of these definitions utilise the xspace package, so the correct amount of space will be automatically inserted after each command.

A University Regulations

The following information has been extracted and paraphrased from the University regulations with regards to formating theses (as presented in the publication: “Post Graduate Study Programme 1994/5”). The csthesis class has been developed with close attention to these regulations.

Paper Size : Thesis should be presented on A4 paper, with all text produced entirely on rectos (single-sided).

Margins : 4cm binding margin, 2cm head margin, 2.5cm fore-edge margin and 4cm tail margin. Page numbers can be countained within the tail margin. Headers should be avoided.

Spacing : Where typescript is used, the main text should be at least 1.5 spacing (18 points leading). Quotations and footnotes should be single spaced.

Character Size : Character sizes should be at least 10 points throughout. The body text (text other than headings) should not exceed 12 points.

Character Font : A serif font should be used.

Word Spacing : There should be an even spacing between words for any particular line. Hyphenation should be avoided if possible.

Title Page : should contain: title of thesis, author’s name, (at the foot of the page) the name of the degree, “University of Edinburgh”, year of presentation.

Bibliography : The bibliography may be single spaced.

B Two Sided Printing

As well as submitting a copy of your thesis to the University, you will also submit a copy to the department. This copy is stored in the Reports and Theses room (1417). Currently the department provide these in an A5 booklet format. You are therefore required to print your thesis using the -Zbook option of lpr for the departmental copy.

You should also specify the twoside class option to your document and recompile the L^{A}TEX source with this option. This will ensure that margins are correct for two-sided printing and that blank pages are inserted where required (e.g. directly after the title page, after the dedication page, and at the end of the document).

C Example L^{A}TEX 2ε Code

%
% A short example of how to use the csthesis class to
% write a thesis which corresponds to university
% regulations. Version 1.1. M. Reddy (mxr), 13 May 1996.
%
\documentclass[phd]{csthesis}
\begin{document}
\title{How I Did It}
\author{Victor Frankenstein}
\submityear{1785}
\maketitle
\dedication{To my Mum.}
\begin{abstract}
This doctoral thesis will present the results of my
work into the reanimation of lifeless human tissue.
\end{abstract}
\begin{acknowledgements}
Many thanks to my mummy for the numerous packed lunches;
and of course to Igor, my faithful lab assistant.
\end{acknowledgements}
\standarddeclaration
\tableofcontents
\chapter{Experimental Procedure}
First you have to get some dead tissue (the knack here is in also
managing to evade a short custodial sentence for grave-robbing).
Then you need one of those crackling things which makes lots of
sparks. The work must also be conducted in a suitably dark and
ancient castle, in close proximity to a town of highly superstitious
and pitch-fork-wielding peasants.
The remaining experimental procedure and theoretical mumbo-jumbo is
left as an excercise for the interested reader.
\end{document}
     Above L^{A}TEX source available on-line in /usr/local/share/TeX/tex/latex/examples